OpenStudy (anonymous):

Solve by substitution; -5x-7y=9 -2x+y=-23 please/thanks:)

hero (hero):

In general, you want to solve for one variable in one equation, and then put the expression for that variable into the other equation.

hero (hero):

You can do that, or you can solve it this way: -5x-7y=9 -2x+y=-23 y = (9 + 5x)/-7 y = -23 + 2x y = y -7(-23 + 2x) = 9 + 5x 161 - 14x = 9 + 5x 161 - 9 = 14x + 5x 152 = 19x 152/19 = x 8 = x -7 = y

hero (hero):

Okay, I'm about to go grocery shopping now. Catch you later.

OpenStudy (anonymous):

THANK YOU!